Ticket Number: HEA-NOC/20080116-13 Ticket Status: UPDATE
Ticket Type: scheduled Resolver:
Ticket Opened: 20080116 17:53 UTC Problem Start: 20080117 09:30 UTC
Ticket Update: 20080117 10:51 UTC Problem End:
NCI currently connects to the HEAnet network using an ATM interface.
New fibre has been installed which will support Gigabit connectivity
between NCI and the HEAnet CRS-1 core routers (although this will
initially operate at 100mbps).
Two new peerings will be brought up at 09:30am Thursday 17th Feb.
The switchover will occur between 09:30 and 10:00 on Thursday 17th Feb.
No outage is expected but NCI Internet connectivity should be considered
at risk during this window. Some asymmetric BGP routing may occur during
the switch, resulting in latency fluctuaton.
After the work NCI will have diverse connectivity between the edge of
their network and the HEAnet network.
20080116 17:39 PC - Interface IPs have been applied to the NCI router and
the HEAnet CRS-1 core routers. The link has been tested and BGP peerings
will be established at tomorrow morning.
20080117 10:37 PC - The link to cr2-cwt was brought up and traffic from
HEAnet to NCI is flowing over that link. The much lower volume of traffic
from NCI to HEAnet continued to flow over the ATM link, which was not
expected. The new peering has been shutdown to maintain symmetric traffic
This ticket can be monitored at http://www.hea.net/tickets/20080116-13
HEAnet Limited. Registered in Ireland, No. 275301.